OpenStudy (anonymous):

Find the common difference for the given sequence? 56, 49, 42, 35, ...

OpenStudy (amistre64):

have you learned how to subtract yet?

OpenStudy (amistre64):

56-49=k 49-42=k 42-35=k 35-next term = k etc ....